OpenStudy (anonymous):

Find the hypotenuse of a right triangle whose legs are sqrt 10 and sqrt 15. sqrt 5 5 15

OpenStudy (welshfella):

x^2 = (sqrt10)^2 + (sqrt15)^2

OpenStudy (welshfella):

note: by definition (sqrt2)^2 = 2

OpenStudy (anonymous):

15?

OpenStudy (welshfella):

Yes 15 = (sqrt15) ^2

OpenStudy (er.mohd.amir):

h*h=10+15 h*h=25 h=5

OpenStudy (anonymous):

thanks

OpenStudy (welshfella):

yw

OpenStudy (er.mohd.amir):

answer is 5 how it 15?